Sketch the graph of

y=1/π+r2

by finding the domain, symmetries, critical points, inflection points,

intercept points, asymptotes, extremas, intervals on which the function is increasing or decreasing,

concave up or down.




1
Expert's answer
2021-01-20T02:52:10-0500

"y=\frac{1}{\pi+r^2}"


Domain: "r\isin(-\infin,\infin)"


Asymptote:

"\displaystyle\lim_{r\to -\infin}f(r)=\displaystyle\lim_{r\to \infin}f(r)=0"

Horizontal asymptote is "y=0"


Symmetry:

there is symmetry respect to y-axis: "f(r)=f(-r)"


Critical point, extrema:

"y'=-\frac{2r}{(\pi+r^2)^2}=0\implies r=0"

the function is increasing on "r\isin(-\infin,0)" , "y'>0"

the function is decreasing on "r\isin(0,\infin)" , "y'<0"

maxima is "(0,1/\pi)"


Inflection points:

"y''=-\frac{2(\pi+r^2)^2-2r\cdot4r(\pi+r^2)}{(\pi+r^2)^4}=0"

"4r^2-\pi-r^2=0"

"r=\pm\sqrt{\pi/3}"

Inflection points are "(\sqrt{\pi/3},\frac{1}{\pi+3}), (-\sqrt{\pi/3},\frac{1}{\pi+3})"


Intercept point:

"r=0\implies y=1/\pi"

y-intercept point is "(0,1/\pi)"

there no x-intercepts